Opened 8 years ago

Closed 7 years ago

Last modified 14 months ago

#411 closed flight processing (fixed)

EUFAR11/07, flight day 220/2011, Toledo

Reported by: knpa Owned by: knpa
Priority: alpha 4 high Milestone: 2011 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by knpa)

Data location: /users/rsg/arsf/arsf_data/2011/flight_data/spain_portugal/EUFAR11_07-2011_220_Toledo/

Data arrived from ARSF via network transfer on 16/08/2011

EUFAR11_07 Scientific objective:
Soil erosion study.

EUFAR code: SEDMEDHY

Priority: a4h

PI: Thomas Schmid

Sensors:

Camera (18/11/11)
Eagle (18/11/11)
Hawk (18/11/11)
LiDAR (18/11/11)

Change History (37)

comment:1 Changed 8 years ago by knpa

  • Summary changed from EUFAR11_07, flight day 220/2011, Toledo to EUFAR11/07, flight day 220/2011, Toledo

comment:2 Changed 8 years ago by knpa

We are currently awaiting Basestation data from Thomas Schmid.

comment:3 Changed 8 years ago by knpa

We have received the data referenced above.

comment:4 Changed 8 years ago by knpa

  • Owner set to knpa
  • Status changed from new to assigned

Beginning LiDAR processing.

comment:5 Changed 8 years ago by knpa

Completed IPAS processing.

Unfortunately result was is less than satisfactory for the first flightline, where the arsf basestation loses it's fix. This could be not improved upon by adding the permanent basestations. The 1-second data from the permanent stations are unavailable for this day apart for SONS (which is a little way away).

comment:6 Changed 8 years ago by knpa

Slight horizontal discrepancies which looked like it was caused by a pitch error.
Changed pitch from -0.00324419 to -0.003 which fixed the problem.

comment:7 Changed 8 years ago by knpa

This flight suffers from roll error. It is not particularly severe though: there will be a vertical error of approximately 10cm at the edge of the swath width.

The roll discrepancy is not constant between flightlines so will not attempt to correct. Unsure if the error is random. Early flightlines have a positive roll discrepancy but this then seems to switch to negative by the fourth flightline.

comment:8 Changed 8 years ago by knpa

The first cross line (14:33) has a large discrepancy with other lines - approximately 45cm.

comment:9 Changed 8 years ago by knpa

There is a lot of noise for this flight.

comment:10 Changed 8 years ago by knpa

Finished QC, preparing delivery.

comment:11 Changed 8 years ago by knpa

  • Description modified (diff)

comment:12 Changed 8 years ago by knpa

LiDAR ready for delivery check:

EUFAR11_07-220-lidar-20110921

comment:13 Changed 7 years ago by knpa

Eagle line 1 is missing dark lines.

comment:14 Changed 7 years ago by knpa

Eagle line 1 number of lines in header was incorrect, have corrected from 6351 to 6352.

comment:15 Changed 7 years ago by jaho

LiDAR delivery check:

  • No README_DEM_MAY_NOT_BE_ACCURATE.txt in dem directory.
  • Description in dem hdr files doesn't match the names of the files.
  • There seems to be an issue similar to that with LiDAR data from Italy. The flights have been flown at altitude of over 5000 meters (16000 feets) and the intensity values are very low, mostly below 50 and AGC is 255 for most of the points.

Otherwise ready to go.

comment:16 Changed 7 years ago by jaho

Started camera processing.

comment:17 Changed 7 years ago by jaho

Removed camera images 1 to 9 - taken before the flight.

comment:18 Changed 7 years ago by jaho

Removed camera images 10 to 12 - overexposed.

comment:19 Changed 7 years ago by jaho

Camera ready for delivery check.

comment:20 Changed 7 years ago by jaho

Rechecked LiDAR. Looks excellent. Ready to deliver.

comment:21 Changed 7 years ago by knpa

e h
1 -0.02 -0.07
2 -0.03 -0.02
3 -0.05 -0.06
4 -0.04 -0.06
5 -0.06 -0.02
6 -0.05 -0.05
7 -0.02 -0.06
8 -0.08 -0.04

comment:22 Changed 7 years ago by adbe

RCD delivery check: No problems, ready to deliver

comment:23 Changed 7 years ago by knpa

Hyperspectral complete and ready for delivery check at 'EUFAR11_07-220-hyperspectral-20111116'

comment:24 follow-up: Changed 7 years ago by jaho

Checked hyperspectral.

An outstanding delivery except few minor issues:

  • Incorrect DEM name in hdr file
  • The name of DEM file in the ReadMe for example aplcorr command doesn't match the name of the DEM in dem/ directory.
  • No mention of blank first and last bands in eagle files under quality remarks in the ReadMe
  • There seems to be some variation of the SCT values within the flightlines between different bands(?). Eagle bands: 154-156, hawk bands: 68-74 and 153-168.

comment:25 in reply to: ↑ 24 Changed 7 years ago by benj

Replying to jaho:

  • There seems to be some variation of the SCT values within the flightlines between different bands(?). Eagle bands: 154-156, hawk bands: 68-74 and 153-168.

We had this issue once before (can't find the reference now) - after investigation we believe it to be an artifact of how fastqc displays bands with different relative signal levels, not a problem with the actual data.

comment:26 Changed 7 years ago by knpa

Fixed the other minor issues mentioned above. Putting on disk to be delivered with LiDAR and RCD.

comment:27 Changed 7 years ago by knpa

  • Component changed from Processing: general to Archiving
  • Description modified (diff)

All sensors delivered today (18/11/11) to Thomas Schmid at Ciemat in Spain.

comment:28 Changed 7 years ago by emca

PI visited ARSF 1-2nd Dec and took a copy of the LiDAR, Camera and Hyperspectral deliveries.

We also re-created an expanded LiDAR/ASTER DEM as the one within the delivery did not completely cover the flightlines for apl processing.

comment:29 Changed 7 years ago by mark1

Hyperspectral hawk bad pixel files and XML files have been re-created and re-delivered to PI (via email). This corrects sample number in bad pixel text files and Eagle IFOV value in XML and extra comments included.

comment:30 Changed 7 years ago by knpa

Beginning archiving.

comment:31 Changed 7 years ago by knpa

Prepared for archiving. Will make available to NEODC once we've sorted out our new procedure.

comment:32 Changed 7 years ago by knpa

Archived by NEODC, moved to archive disk.

comment:33 Changed 7 years ago by knpa

  • Status changed from assigned to accepted

comment:34 Changed 7 years ago by knpa

  • Resolution set to fixed
  • Status changed from accepted to closed

comment:35 Changed 15 months ago by dac

Mapped hyperspectral data and created new delivery by hardlinking old one and adding mapped files to it.

comment:36 Changed 14 months ago by asm

Archiving

Mapped files were not present in the hyperspectral delivery. Created the mapped files and a new delivery in order to create as well JSON files for the EUFAR flight finder tool. New delivery and JSON files are being uploaded to CEDA (started on 15/01/2018).

comment:37 Changed 14 months ago by asm

Archiving

Upload finished, CEDA notified on 18/01/2017.

Note: See TracTickets for help on using tickets.